9. In 2016, Olive’s Optometry, a sole proprietorship, purchased a piece of business-use property (Equipment A) for $1,000,000. Olive’s owned Equipment A for more than a year and depreciation deductions totaled $600,000.

Assume Olive’s sold Equipment A for $1,150,000 in 2019, and Olive’s paid $50,000 in selling expenses. What is Olive’s recognized gain on the sale of Equipment A in 2019?

10.What amount of Olive’s recognized gain is recaptured as ordinary income (1245 recapture)?

11.  What is the amount of Olive’s 1231 gain on the sale of Equipment A?

12. In 2016, Olivia’s Optometry, a sole proprietorship, purchased a piece of business-use property (Equipment A) for $1,000,000. Olivia’s owned Equipment A for more than a year and depreciation deductions totaled $600,000.In 2019, Olivia’s sold Equipment A for $200,000, and Olivia’s paid no selling expenses. What amount of Olivia's recognized gain or loss is recaptured as ordinary income under section 1245?
Answer #1
A Cost of Equipment $1,000,000
B Depreciation deduction $600,000
C=A-B Book Value at the time of sale $400,000
D Selling price $1,150,000
E Selling expenses $50,000
F=D-E Net Sales receipt $1,100,000
G=F-C Recognized gain on sale $700,000
H=A-C 1245 recapture as ordinary income $600,000
I=F-A 1231 gain (Capital gain) $100,000
12
A Cost of Equipment $1,000,000
B Depreciation deduction $600,000
C=A-B Book Value at the time of sale $400,000
D Selling price $200,000
E=C-D Recognized loss $200,000